State whether the following sequences are arithmetic; and if so, state the common difference.
15, 18, 21, 24

Answer by Edwin McCravy(20056)

15, 18, 21, 24

How to tell if it is arithmetic.

1. Subtract the 1st term from the 2nd term.  Notice what you get. 18-15 = 3
2. Subtract the 2nd term from the 3rd term.  Notice what you get. 21-18 = 3
3. Subtract the 3rd term from the 4th term.  Notice what you get. 24-21 = 3

Since we got the same difference 3 each time we subtracted a term from the
next one, the difference 3 was in common.  So we call 3 a "common difference"
and the sequence is an arithmetic sequence.

Edwin